OpenStudy (anonymous):

How would you solve for n? n!/(n-2)! = 42

jimthompson5910 (jim_thompson5910):

by definition n! = n*(n-1)*(n-2)*...*3*2*1 we can rewrite that to say n! = n*(n-1)*(n-2)!

jimthompson5910 (jim_thompson5910):

so \[\Large \frac{n!}{(n-2)!} = 42\] \[\Large \frac{n(n-1)(n-2)!}{(n-2)!} = 42\] \[\Large \frac{n(n-1)\cancel{(n-2)!}}{\cancel{(n-2)!}} = 42\] \[\Large n(n-1) = 42\] I'll let you finish

OpenStudy (anonymous):

ok thank you

jimthompson5910 (jim_thompson5910):

np